OpenStudy (anonymous):

Which of the following shows the equation below correctly solved for a? b = 4a − 17 − 2b A. a=3/4b+17 B. a=3/4b-17 C. a=3b-17/4 D. a=3b+17/4

OpenStudy (anonymous):

are you sure it isn't y =4a − 17 − 2b

OpenStudy (anonymous):

no

OpenStudy (anonymous):

im going with D I got 3b/4 + 17/4 for a

OpenStudy (anonymous):

the closest thoguh

OpenStudy (anonymous):

If not, then you solve by combining like terms. 3b + 17 / 4 = a

OpenStudy (anonymous):

That's the right answer: D

OpenStudy (anonymous):

b = 4a − 17 − 2b +2b ----------------- 3b = 4a -17 +17 ------------- 3b + 17 = 4a ------- --- 4 4 3b +17/4 =a